The sum of two numbers is 111 and their difference is 63. What are the two numbers?
a + b = 111
a = 111-b
a - b = 63
I can't figure out how to solve.

Answer by checkley71(8403)

A+B=111
A-B=63
THEREFORE A=B+63 SUBSTITUTING A=B+63 IN THE FIRST EQUATION WE GET
B+63+B=111 OR 2B+63=111 OR 2B=111-63 OR 2B=48 THEREFORE B=24 & A+24=111 OR
A=111-24 OR A=87 AND 87-24=63.